# LLM Output Homogenization is Task Dependent

### v1

@@ 4.1 Experiment Details
Models    We use three generative models: GPT-4o, Claude-4-Sonnet, and Gemini-2.5-Flash. These models are used in our evaluation to generate responses and separately as judges of task-anchored functional diversity (temperature 0). When reporting LLM-judge metrics, we average the outputs across these three models.
@@
Datasets    We evaluate $n=300$ prompts from a variety of datasets that achieve reasonable coverage [^2] We did not identify any datasets containing prompts for category E (Problem Solving or Design Subjective), likely due to the inherently open-ended nature of such tasks. We include it in our taxonomy for symmetry with category D. Developing and open-sourcing datasets in this category may be a valuable direction for future work. across the task categories in our taxonomy (see Appendix Table 3 for counts). The five datasets used in our evaluation are: Community Alignment (Zhang et al., 2025a), MATH-500 (Lightman et al., 2023), NoveltyBench (Zhang et al., 2025b), SimpleQA (Wei et al., 2024), and WildBench (Lin et al., 2025). These datasets represent a mix of user-generated and curated prompts. Appendix 6.1 includes additional details about each dataset and how we sampled prompts.
Sampling Strategies    To evaluate homogenization over multiple responses, we compare three sampling strategies: temperature sampling, system prompt sampling, and in-context regeneration. For each sampling strategy, we sample $5$ responses per prompt. For temperature sampling, we consider three temperature levels for each model based on its permitted range: low ($t=0.0$), medium ($t=1.0$ for GPT and Gemini, $t=0.5$ for Claude), and high ($t=2.0$ for GPT and Gemini, $t=1.0$ for Claude). Recall that system prompt sampling refers to using a system prompt that instructs the model to generate a specified number of responses. The model is instructed to produce multiple responses separated by a delimiter, allowing them to be de-aggregated with regular expressions. In-context regeneration samples responses by iteratively prompting the model to generate a different response, while keeping all previous responses in context. The first response is generated with temperature sampling. For both system prompt sampling and in-context regeneration, we use the medium temperature values.
@@
We further evaluate both general and task-anchored approaches to system prompt sampling and in-context regeneration. For the general approach, we use a variation of the prompts used previous works (Zhang et al., 2025a, b). Our task-anchored approach modifies these prompts to specify the functional difference relevant to each task category in our taxonomy. Table 2 shows how we modify each instruction to be task-anchored for two task categories. Appendix 6.3 provides all our task-anchored prompts for system prompt sampling and in-context regeneration.
@@
Diversity Metrics    We compute three diversity metrics: task-anchored functional diversity (Def. 3.1), vocabulary diversity (Def. 3.2), and embedding diversity (Def. 3.3). To calculate functional diversity, we use task-anchored LLM-judges [^3] We validate the LLM-judges on a stratified random sample of 225 response pairs across models, tasks, and sampling strategies. Two authors independently labeled these responses for functional diversity, and agreed 79% with the LLM-judges (80% agreement between annotators). This is comparable to Zhang et al. (2025b)’s evaluation of functional diversity in creative tasks, where a fine-tuned judge also achieved 79% agreement with human labels. Appendix Table 5 reports detailed inter-rater agreement. , where the judge prompt includes the definition of functional difference for the relevant task category (see Appendix 6.4 for judge prompts). We then use these pairwise comparisons [^4] Pairwise comparisons grow quadratically, which is why we only generate 5 responses per prompt. to determine the number of functionally diverse responses obtained (out of the 5 responses generated per prompt and sampling strategy). Specifically, we group responses that are judged as not functionally different from each other, and then count the number of unique groups (connected components). To compute embedding diversity, we generate response embeddings using the gemini-embedding-001 model (with 3072-dimensional embeddings).
@@
Quality Metrics    We evaluate the quality of responses in two ways. (1) Reward Model Quality: Following many recent work that evaluates the diversity-quality trade-off (Lanchantin et al., 2025a; Slocum et al., 2025), we measure quality in terms of reward scores assigned by a reward model. We use Athene-RM-8B, which is to date empirically validated as one of the best reward models for human preferences (Frick et al., 2025). (2) Checklist-Based Quality: We also measure quality following prior work that uses LLM-judges with grading checklists (Lin et al., 2025; Wei et al., 2025). In this approach, the LLM-judge first generates a checklist of 3-5 key factors for response quality in a given prompt. This prompt-specific checklist is then used by the LLM-judge to score a particular response on a Likert scale from 1 to 5, where 1 indicates that none of the checklist criteria are met and 5 indicates that all criteria are satisfied. We include the judge prompts and examples of generated checklists in Appendix 6.5.
@@
4.2 Functional Diversity
We report our evaluation results on functional diversity. Our main finding is that our task-anchored sampling technique outperforms the more general sampling techniques in previous work (Zhang et al., 2025a, b). Figure 2 shows how we significantly increase functional diversity for task categories where homogenization is undesired, while preserving homogenization where it is desired (for GPT-4o, all results in Appendix 7). Below, we explore results across task categories.
Well-Specified Objective Tasks (Category A)    Tasks in this category have a single verifiably correct answer; therefore, no functional diversity is expected. But when employing general diversity-promoting sampling methods, homogenization is undesirably reduced, as evidenced by the generation of multiple unique answers ($2$ on average). Increasing temperature also undesirably reduces homogenization. In contrast, our task-anchored sampling method maintains homogenization, consistently producing one unique answer per task.
Underspecified Objective and Random Generation Tasks (Categories B & C)    Tasks in these categories are characterized by the existence of multiple verifiably correct answers, which suggests that models may easily conceptualize what difference means here. Consequently, we observe no significant differences between task-anchored and general sampling approaches, as the concept of diversity—defined as producing distinct correct answers—is inherently straightforward in this context. Both methods yield nearly maximal functional diversity, with approximately 5 unique responses out of 5 generations. In contrast, higher temperature settings result in suboptimal functional diversity, producing only 2 to 3 unique responses on average.
Problem-Solving Objective Tasks (Category D)    Tasks in this category are defined by the presence of a single correct answer, but allow for multiple valid explanations or solution strategies. In this setting, we find that neither temperature-based sampling nor general strategies are effective in eliciting responses with diverse solution strategies. In contrast, both task-anchored system prompts and in-context regeneration sampling are able to generate approximately 2–3 distinct solution strategies. This relatively low number may be attributable to the inherent difficulty of the MATH-500 benchmark, which poses significant challenges for large language models in producing even a single correct solution (Hendrycks et al., 2021).
Partial and Non-Verifiable Tasks (Categories F, G, H)    Tasks in these categories cover encyclopedia inquiries, creative writing, and requests for advice or opinions. Across all three models, our task-anchored sampling methods – both system prompt and in-context regeneration – significantly reduce homogenization compared to their respective general approaches (t-test, $p<0.05$). For GPT-4o and Gemini-2.5-Flash, task-anchored system prompting yields the highest number of functionally diverse responses. In contrast, for Claude-4-Sonnet, both task-anchored methods demonstrate comparable performance in promoting response diversity.
@@ 4.3 Diversity-Quality Tradeoff: Comparing General & Task-Based Metrics
We find that improved functional diversity from our task-anchored sampling does not decrease the quality of responses, when the task-dependent nature of quality is captured in the quality metric. [^5] For tasks with singular verifiable rewards (Simple-QA & MATH-500), we separately validate the accuracy of responses (Appendix Table 13). Overall, our task-anchored sampling approaches often improve accuracy. For Simple-QA, our task-anchored sampling outperforms temperature sampling for all 3 models. For MATH-500, our task-anchored sampling outperforms temperature sampling for Claude-4-Sonnet and Gemini-2.5-Flash, and maintains accuracy for GPT-4o. Recent proposals for measuring quality using task-specific checklists align with our discussions around task-anchored metrics for diversity (Lin et al., 2025; Wei et al., 2025). Whereas, when quality is determined by a reward model, the scores do not inherently reflect task differences (e.g. the quality of a creative writing response is measured in the same way as the quality of a math problem-solving response). Figure 3 shows that the diversity-quality tradeoff prevalent in previous studies may simply be the result of mis-conceptualizing both diversity and quality. When evaluating general metrics, there appears to be a large diversity-quality tradeoff between vocabulary diversity and reward quality (Figure 3(a)-3(c)), and the tradeoff is similarly large with embedding diversity (Appendix Figure 6(c)). When we compare task-anchored functional diversity with checklist-based quality, there is a negligible diversity-quality tradeoff (Figure 3(d)-3(f)). In particular, our task-anchored sampling technique maintains the same level of quality as the general strategies in previous work (Zhang et al., 2025a, b), while improving functional diversity.
(a) General Metrics: GPT-4o
(b) General Metrics: Claude-4-Sonnet
(c) General Metrics: Gemini-2.5-Flash
(d) Task-Based Metrics: GPT-4o
(e) Task-Based Metrics:
Claude-4-Sonnet
(f) Task-Based Metrics:
Gemini-2.5-Flash
@@
Figure 3: With task-based metrics, diversity is improved with no significant drop in quality. We plot quality on the $x$-axis and diversity on the $y$-axis and compare the tradeoff under general metrics vs task-based metrics. In (a)-(c), there is a large tradeoff between vocabulary diversity (Def. 3.2) and quality scores determined by a reward model. In (d)-(f), there is a negligible tradeoff between task-anchored functional diversity (Def. 3.1) and LLM-judges with task-based grading checklists. Note that the checklist-based quality difference between score 44 and $5$ is “good” vs “very good”. Plots show the mean and standard error of all metrics averaged across all task categories except category A, which we exclude because it is the only category where output homogenization is desired.
@@ 5 Discussion
Our work underlines the task-dependent nature of evaluating and mitigating output homogenization. We find that our task-anchored sampling technique outperforms more general sampling approaches in terms of increasing response diversity only when desired. Our results show that without task-dependence, previous methods to analyze output homogenization often (1) misconceptualize output diversity (2) reduce homogenization in tasks where homogenization should be preserved and (3) maintain homogenization in tasks where more pluralism is desired. Further, our results show that task-anchored sampling does not result in a significant diversity-quality trade-off. These results challenge the perceived existence of a diversity-quality tradeoff that is common in the literature. In this section, we discuss the implications of our work and avenues for future research.
@@ 5.1 Our framework improves homogenization evaluation
We have developed a taxonomy of task categories that clarifies how a model can conceptualize diversity based on the categorization. For example, evaluating homogenization in math problem-solving should measure variety in solution strategies, whereas evaluating homogenization in advice or opinions should measure variety in viewpoints or perspectives. We improve upon previous studies that rely on generic measures of diversity (vocabulary or embedding differences), which is particularly meaningful when evaluating diversity-promoting methods. Our findings suggest that using general metrics without accounting for task dependence does not capture meaningful functional diversity.
We highlight the importance of evaluating prompts across our taxonomy when analyzing output homogenization. When studies limit their evaluation to tasks where diversity is desired, there may be unintended effects (e.g. confabulations) when those methods are applied to tasks which rely on homogenization being preserved. Hence, not adopting a task-dependent approach could result in less robust evaluation and present safety or ethical concerns downstream. Our taxonomy is one example of a categorization that anchors task dependence. Future work can adapt or expand this categorization. Our framework is generalizable in that the task-anchored approach can support a custom taxonomy tailored to a more specific downstream use case.
@@ 5.2 Our framework improves homogenization mitigation
There are many ways to apply task-dependence in practical settings and our approach could be applied at inference-time automatically. For instance, the model could be given (or instructed to determine) a task categorization, filter prompts into said categories, and output responses according to the task-based conceptualization of output homogenization. Our main improvement is in clarifying model instructions for output homogenization behavior in terms of the task category. Instead of assuming the model does this inherently, it may be important to clarify and steer expected behavior.
Although we focus on prompt-based strategies, our task-anchored approach may be applied to other diversity-promoting methods that modify the alignment process. For example, Lanchantin et al. (2025a) propose a method for improving diversity through preference pair construction $(x,y^{+},y^{-})$. This approach could be modified to construct pairs in a task-informed way that avoids learning undesired semantic preferences that might reduce functional diversity. Slocum et al. (2025) also propose modifying the RLHF or DPO optimization objective to include a penalty for lower token-level entropy. This penalty could be selectively applied to certain task categories where vocabulary diversity is desired, such as random generation and creative writing.
@@
  • 6.1: Evaluation Datasets
  • 6.2: Task Classification Into Our Taxonomy
  • 6.3: Prompts for Task-Anchored Sampling Strategies
  • 6.4: Measuring Functional Diversity
  • 6.5: Measuring Quality Using LLM-Judges With Task Grading Checklists
@@
  • Figure 4: Functional Diversity for Claude-4-Sonnet (c.f. Figure 2 in main text).
  • Figure 5: Functional Diversity for Gemini-2.5-Flash (c.f. Figure 2 in main text).
  • Figure 6(c): Diversity-Quality Tradeoff Using Embeddings (c.f. Figure 3 in main text).
  • Table 8: Functional Diversity by Model, Sampling Strategy, and Task Category
  • Table 9: Vocabulary Diversity by Model, Sampling Strategy, and Task Category
  • Table 10: Embedding Diversity by Model, Sampling Strategy, and Task Category
  • Table 11: Checklist-Based Quality by Model, Sampling Strategy, and Task Category
  • Table 12: Athene-RM-8B Reward by Model, Sampling Strategy, and Task Category
  • Table 13: Accuracy by Model, Sampling Strategy, and Task Category (for verifiable tasks)

### v2

@@ 4.1 Experiment Details
We evaluate responses from five models: GPT-4o, Claude-4-Sonnet, Gemini-2.5-Flash, Llama-3.1-8B-Instruct, and Mistral-7B-Instruct-v0.3. Separately, we use GPT-4o, Claude-4-Sonnet, and Gemini-2.5-Flash as LLM-judges (independent of response generation). When reporting LLM-judge metrics, we average the outputs across these three judge models.
Datasets
@@
We evaluate $n=344$ prompts from a variety of datasets that achieve reasonable coverage across the task categories in our taxonomy (c.f. Appendix Table 6). The 6 datasets used in our evaluation are: Community Alignment (zhang2025cultivating), MacGyver tian2024macgyver, MATH-500 (lightman2023let), NoveltyBench (zhang2025noveltybench), SimpleQA (wei2024simpleQA), and WildBench (linwildbench). These datasets represent a mix of user-generated and curated prompts. Appendix 6.2 includes more details about each dataset and how we sampled prompts.
Sampling Strategies
@@
To evaluate homogenization over multiple responses, we compare three sampling strategies: temperature sampling, system prompt sampling, and in-context regeneration. For each sampling strategy, we sample $5$ responses per prompt. For temperature sampling, we consider three temperature levels for each model based on its permitted range: low ($t=0.0$ for GPT, Claude, and Gemini, $t=0.1$ for Llama and Mistral), medium ($t=1.0$ for GPT and Gemini, $t=0.5$ for Claude, Llama, and Mistral), and high ($t=2.0$ for GPT and Gemini, $t=1.0$ for Claude, Llama and Mistral). We further evaluate both general and task-anchored approaches to system prompt sampling and in-context regeneration (§ 3.3). For the general approach, we use a variation of the prompts used previous works (zhang2025cultivating; zhang2025noveltybench). Our task-anchored approach modifies these prompts to specify the functional difference relevant to each task category in our taxonomy, based on the model’s self-categorization of the task. Appendix 6.4 provides all our task-anchored prompts for system prompt sampling and in-context regeneration. For both system prompt sampling and in-context regeneration, we use the medium temperature values.
Diversity Metrics
@@
We compute four diversity metrics: task-anchored functional diversity (Def. 3.1), vocabulary diversity (Def. 6.1), embedding diversity (Def. 6.2), and compression diversity (Def. 6.3). To calculate functional diversity, we use LLM-judges [^2] We validate the LLM-judges on a stratified random sample of 225 response pairs across models, task categories, and sampling strategies. Two authors independently labeled these responses for functional diversity, and agreed 79% with the LLM-judges (80% agreement between annotators). See Appendix Table 8 for details., where the judge prompt includes the functional diversity concept for the ground-truth task category (see Appendix 6.5 for judge prompts). We determine the ground-truth task category for each prompt based on the source dataset and human annotation (Appendix 6.3). We then determine the number of functionally diverse responses (Def 3.2) out of the 5 responses [^3] Pairwise comparisons grow quadratically, which is why we only generate 5 responses per prompt. generated per prompt and sampling strategy. To compute embedding diversity, we generate response embeddings using the gemini-embedding-001 model. To compute compression diversity, we use Gzip following the method in shaib2024standardizing.
Quality Metrics
@@
We evaluate the quality of responses in two ways. (1) Reward Model Quality: Following many recent work that evaluates the diversity-quality trade-off (lanchantin2025diverse; slocumdiverse), we measure quality in terms of reward scores assigned by a reward model. We use Athene-RM-8B, which is to date empirically validated as one of the best reward models for human preferences (frick2025athene). (2) Checklist-Based Quality: We also measure quality following prior work that uses LLM-judges with grading checklists (linwildbench; wei2025rocketeval). In this approach, the LLM-judge first generates a checklist of 3-5 key factors for response quality in a given prompt. This prompt-specific checklist is then used by the LLM-judge to score a particular response on a Likert scale from 1 to 5, where 1 indicates that none of the checklist criteria are met and 5 indicates that all criteria are satisfied. We manually review all generated checklists, and include the judge prompts and examples of generated checklists in Appendix 6.6.
4.2 Functional Diversity
We first report our evaluation results on functional diversity. Our main finding is that our task-anchored sampling technique outperforms the more general sampling techniques in previous work (zhang2025cultivating; zhang2025noveltybench). Figure 2 shows how we significantly increase functional diversity for task categories where homogenization is undesired, while preserving homogenization where it is desired (for GPT-4o, all results in Appendix 7). Below, we explore results across task categories.
@@ 4.2 Functional Diversity
Well-Specified Objective Tasks (Category A)
Tasks in this category have a single verifiably correct answer; thus, no functional diversity is expected. However, when employing general diversity-promoting sampling methods, homogenization is undesirably reduced, as evidenced by the generation of multiple unique answers ($2$ on average). Increasing temperature also undesirably reduces homogenization. In contrast, our task-anchored sampling method maintains homogenization, consistently producing one unique answer per task for GPT-4o, Gemini-2.5-Flash, and Claude-4-Sonnet.
Underspecified Objective and Random Generation Tasks (Categories B & C)
Tasks in these categories are characterized by the existence of multiple verifiably correct answers, which suggests that models may easily conceptualize what difference means here. Consequently, we observe no significant differences between task-anchored and general sampling approaches, as the concept of diversity—defined as producing distinct correct answers—is inherently straightforward in this context. Both methods yield nearly maximal functional diversity, with approximately 5 unique responses out of 5 generations. In contrast, higher temperature settings result in suboptimal functional diversity, producing only 2 to 3 unique responses on average.
Problem-Solving Objective Tasks (Category D)
Tasks in this category are defined by the presence of a single correct answer, but allow for multiple valid explanations or solution strategies. In this setting, we find that general prompt-based strategies are not effective in eliciting responses with diverse solution strategies. In contrast, both task-anchored system prompts and in-context regeneration sampling are able to generate approximately 2–3 distinct solution strategies. This relatively low number may be attributable to the inherent difficulty of the MATH-500 benchmark, which poses significant challenges for LLMs in producing even a single correct solution (hendrycks2021measuring).
Partial and Non-Verifiable Tasks (Categories E, F, G, H)
Tasks in these categories cover subjective problem-solving, encyclopedia inquiries, creative writing, and requests for advice or opinions. Across all five models, our task-anchored sampling methods – both system prompt and in-context regeneration – reduce homogenization compared to their respective general approaches). For GPT-4o, Gemini-2.5-Flash, and Mistral-7B-Instruct-v0.3, task-anchored system prompting yields the highest number of functionally diverse responses. In contrast, for Claude-4-Sonnet and Llama-3.1-8B-Instruct, both task-anchored methods demonstrate comparable performance in promoting response diversity. Among temperature-sampled outputs, smaller open-weight models tend to have less homogenization than larger commercial models, possibly due to their less extensive alignment.
@@ 4.3 Diversity-Quality Tradeoff: Comparing General & Task-Based Metrics
We find that improved functional diversity from our task-anchored sampling often maintains the quality of responses, when the task-dependent nature of quality is captured in the quality metric. [^4] For tasks with singular verifiable rewards (Simple-QA & MATH-500), we separately validate the accuracy of responses (Appendix Table 26-27). Overall, our task-anchored sampling approaches often maintain and sometimes improve accuracy compared to temperature sampling. For MATH-500, system prompt sampling has the best accuracy for all models except Mistral. For Simple-QA, in-context regeneration performs the best for Gemini and Mistral, while system prompt sampling performs the best for Claude and Llama. Recent proposals for measuring quality using task-specific checklists align with our discussions around task-based metrics for diversity (linwildbench; wei2025rocketeval). Whereas, when quality is determined by a reward model, the scores do not inherently reflect task differences (e.g. the quality of a creative writing response is measured in the same way as the quality of a math problem-solving response). For GPT-4o responses, Figure 3 shows that the diversity-quality tradeoff prevalent in previous studies may simply be the result of mis-conceptualizing both diversity and quality. When evaluating general metrics, there appears to be a large diversity-quality tradeoff between vocabulary diversity and reward quality (Figure 3(a)), and the tradeoff is similarly large with embedding diversity (Appendix Figure 13). When we compare task-anchored functional diversity with checklist-based quality, there is a negligible diversity-quality tradeoff (Figure 3(b)). These results are similar for Claude-4-Sonnet and Gemini-2.5-Flash (Appendix Figures 9-10).
(a) General Metrics (GPT-4o Responses)
(b) Task-Based Metrics (GPT-4o Responses)
@@
Figure 3: With task-based metrics, diversity is improved with no significant drop in quality. We plot quality on the $x$-axis and diversity on the $y$-axis and compare the tradeoff under general metrics vs task-based metrics. In (a), there is a large tradeoff between vocabulary diversity (Def. 6.1) and quality scores determined by a reward model. In (b), there is a negligible tradeoff between task-anchored functional diversity (Def. 3.1) and LLM-judges with task-based grading checklists. Note that the checklist-based quality difference between score 44 and $5$ is “good” vs “very good”. Plots show the mean and standard error of all metrics averaged across all task categories except category A, which we exclude because it is the only category where output homogenization is desired.
With task-based metrics, the diversity-quality tradeoff is more noticeable for smaller open-weight models (Appendix Figures 11-12), but still small ($\sim$0.5 on our 5-point scale). One reason may be that smaller open-weight models have much lower task classification accuracy (only about 50% compared to 85% for commercial models). For commercial models, the quality slightly decreases when generating more than 5 responses using prompt-based methods (Appendix Figure 14). In particular, system prompt sampling may have lower response quality as the number of generated responses approaches the maximum output length for a single generation. Overall, our task-anchored sampling maintains the same level of quality as the general prompt-based strategies in previous work (zhang2025cultivating; zhang2025noveltybench), while improving functional diversity.
@@ 5 Discussion
Our work underlines the task-dependent nature of evaluating and mitigating output homogenization. We find that our task-anchored sampling technique outperforms more general sampling approaches in terms of increasing response diversity only when desired. Our results show that without task-dependence, previous methods to reduce output homogenization often (1) misconceptualize output diversity (2) reduce homogenization in tasks where homogenization should be preserved and (3) maintain homogenization in tasks where more pluralism is desired. Further, our results show that task-anchored sampling does not result in a significant diversity-quality trade-off under task-based metrics. These results challenge the common assumption in the literature of a diversity-quality tradeoff. In this section, we discuss the implications of our work and avenues for future research.
@@ 5.1 Our framework improves homogenization evaluation
We have developed a taxonomy of task categories that clarifies how a model can conceptualize diversity based on the categorization. For example, evaluating homogenization in math problem-solving should measure variety in solution strategies, whereas evaluating homogenization in advice or opinions should measure variety in viewpoints or perspectives. We improve upon previous studies that rely on generic measures of diversity (vocabulary or embedding differences), which is particularly meaningful when evaluating diversity loss in alignment and diversity-promoting methods. Our findings suggest that using general metrics without accounting for task dependence does not capture meaningful functional diversity and may falsely show a diversity-quality tradeoff. Future research may further explore how this applies to alignment. While previous studies show that token entropy collapses during alignment (lanchantin2025bridging), our preliminary experiments show that functional diversity does not necessarily collapse (Appendix Figures 15-16).
We highlight the importance of evaluating prompts across our taxonomy when analyzing output homogenization. When studies limit their evaluation to tasks where diversity is desired, there may be unintended effects (e.g. confabulations) when those methods are applied to tasks which rely on homogenization being preserved. Hence, not adopting a task-dependent approach could result in less robust evaluation and present safety or ethical concerns downstream. Our taxonomy is one example of a categorization that anchors task dependence. An important limitation is that our taxonomy is English-centric; we only define functional diversity concepts in English, and we only evaluate English prompts. Future work may adapt or expand our taxonomy and evaluation approach. To modify the taxonomy and run new evaluations, one simply needs to edit or add new task-anchored prompts for classification, sampling, and evaluation.
Our evaluation relies on LLM-judges to measure task-based diversity and quality, which has known limitations (shi2024judging; li2025generation). Future work could include a user study to confirm that task-based functional diversity aligns with human judgments of what constitutes a meaningful difference between responses. Similarly, future work may further explore how task-dependent quality metrics align with human preferences (wei2025rocketeval; linwildbench).
@@ 5.2 Our framework improves homogenization mitigation
There are many ways to apply task-dependence in mitigating homogenization and our approach could be applied at inference-time automatically. For instance, when given a prompt, the model could be instructed to determine its task categorization and output responses according to the task-based conceptualization of output homogenization. Our main improvement is in clarifying model instructions for output homogenization behavior in terms of the task category. Instead of assuming the model does this inherently, it may be important to clarify and steer expected behavior.
Although we focus on prompt-based strategies, our task-anchored approach may be applied to other diversity-promoting methods that modify the alignment process. For example, lanchantin2025diverse propose a method for improving diversity through preference pair construction $(x,y^{+},y^{-})$ in DPO. This approach could be modified to construct pairs in a task-informed way that avoids learning undesired semantic preferences that might reduce functional diversity. slocumdiverse also propose modifying the RLHF or DPO optimization objective to include a penalty for lower token-level entropy. This penalty could be selectively applied to certain task categories where vocabulary diversity is desired, such as random generation and creative writing. Furthermore, li2025jointly propose Diversity Aware Reinforcement Learning (DARLING) to jointly optimize for response quality and semantic diversity. While they use a general semantic diversity classifier, this approach could be modified to use task-dependent functional diversity. We evaluate and explore modifying DARLING using our task-dependent framework in Appendix 6.7.
@@
  • 6.1: Taxonomy Crosswalks
    • Table 3: Crosswalk of Our Taxonomy and Previous Output Homogenization Studies
    • Table 4: Crosswalk of Task Categories for ChatGPT Usage with Our Taxonomy
    • Table 5: Crosswalk of Task Categories for Claude Usage with Our Taxonomy
  • 6.2: Evaluation Datasets
  • 6.3: Task Classification Into Our Taxonomy
    • Table 6: Number of Prompts per Dataset and Taxonomy Category
    • Figure 4: Recall for Models’ Task Classification with Human Annotation
  • 6.4: Prompts for Task-Anchored Sampling Strategies (Table 7)
  • 6.5: Measuring Functional Diversity
    • Table 8: Agreement Between Human Annotation and LLM-Judges
    • Table 9: Prompts for Functional Diversity LLM-Judges
    • Table 10: Examples of Functionally Diverse Responses by Category
    • Table 11: Examples of Homogeneous Responses by Category
  • 6.6: Measuring Quality Using LLM-Judges With Task Grading Checklists
    • Table 12: Examples of Task-Specific Grading Checklists
    • Table 13: Example Responses Comparing Checklist-Based Grading & Athene Reward
  • 6.7: Alignment Experiments
@@
  • Figure 5, 6, 7, 8: Functional Diversity for Claude-4-Sonnet, Gemini-2.5-Flash, Llama-3.1-8B-Instruct, and Mistral-7B-Instruct-v0.3 (c.f. Figure 2 for GPT-4o)
  • Figure 9, 10, 11, 12: Claude-4-Sonnet, Gemini-2.5-Flash, Llama-3.1-8B-Instruct, and Mistral-7B-Instruct-v0.3 (c.f. Figure 3 for GPT-4o)
  • Figure 13: Diversity-Quality Tradeoff Using Embeddings
  • Figure 14: Diversity-Quality Tradeoff for Varying Number of Generated Responses (5-10)
  • Figures 15-16: Functional Diversity for Llama-3.1-8B-Instruct with DPO and GRPO
  • Figure 17: Diversity-Quality Tradeoff for Llama-3.1-8B-Instruct with DPO and GRPO
  • Figures 18-19: Functional Diversity for Llama-3.1-8B-Instruct with DARLING alignment
  • Figure 20: Diversity-Quality Tradeoff for Llama-3.1-8B-Instruct w/DARLING alignment
  • Table 14-15: Functional Diversity by Model, Sampling Strategy & Task Category
  • Table 16-17: Vocabulary Diversity by Model, Sampling Strategy & Task Category
  • Table 18-19: Embedding Diversity by Model, Sampling Strategy & Task Category
  • Table 22-23: Checklist-Based Quality by Model, Sampling Strategy & Task Category
  • Table 24-25: Athene-RM-8B Reward by Model, Sampling Strategy & Task Category
  • Table 26-27: Accuracy by Model, Sampling Strategy & Task Category (for verifiable tasks)
  • Table 28, 29, 30: Functional Diversity As Calculated by Single LLM-Judges
  • Table 31, 32, 33: Checklist-Based Quality As Calculated by Single LLM-Judges
  • Table 34-35: Functional Diversity & Checklist-Based Quality for 10 Generated Responses
  • Table 36: Functional Diversity for Llama-3.1-8B-Instruct with DPO and GRPO
  • Table 37 Functional Diversity for Llama-3.1-8B-Instruct with DARLING alignment